How do I set this problem up? I tried but I came up with a negative number.
A nurse must mix 15L of a 10% solution of a drug with 60% solution to obtain a 20% mixture. How many liters of the 60% solution will be needed?

Answer by checkley77(12844)

.60X+.10*15=.20(15+X)
.60X+1.5=3+.20X
.60X-.20X=3-1.5
.40X=1.5
X=1.5/.40
X=3.75 LITERS OF 60% SOLUTION IS
PROOF:
.60*3.75+.10*15=.20(15+3.75)
2.25+1.5=.20*18.75
3.75=3.75